What is the present value of $6,000 to be received at the end of 8 years assuming 12% intrest.

A 6,720
B 2,423
C 5,357
D 11,842
Answer #1

Future value = FV = $6000

Time period = n = 8

interest rate = r = 12%

Present value (PV) is calculated using the formula:

PV = FV/(1+r)n

Present Value = PV = 6000/(1+12%)8 = 6000/1.128 = 2423.299368

Answer -> 2,423